# On the Viability of Open-Source Financial Rails: Economic Security of Permissionless Consensus

### v1

@@ 4.1 The Stubborn Nakamoto protocol
The Stubborn Nakamoto Protocol: All consensus nodes run Nakamoto’s blockchain protocol, [^51] We make the standard “implicit echoing” assumption (which is also assumed for standard Nakamoto). That is, every node (including observer nodes that do not mine) echoes every fresh message they see. This implies that if an honest node sees a message at time $t$, then every honest node will have seen it by $t+\Delta$. except for the following changes:
  • Finality: Suppose that $i$ confirms block ${\sf B}$ at time $t$. Node $i$ then waits for $2\Delta$ time. If by time $t+2\Delta$ node $i$ has not confirmed any block ${\sf B}^{\prime}\neq{\sf B}$ that conflicts with ${\sf B}$, [^52] That is, according to the local view of node $i$, at time $t+2\Delta$ block ${\sf B}$ is the only block of that height that is at least $k$-deep in the longest chain. then node $i$ finalizes block ${\sf B}$
  • Halting: Suppose that $i$ confirms block ${\sf B}$ at time $t$, and $i$ confirms block ${\sf B}^{\prime}\neq{\sf B}$ at some time [^53] That is, according to the local view of node $i$, there are two conflicting blocks ${\sf B}\neq{\sf B}^{\prime}$ of the same height, both were at least $k$-deep in the longest chain, and both were confirmed within $4\Delta$ time. $\tau\leq t+4\Delta$. Then node $i$ halts.
  • Ignoring conflicts with finalized blocks: Suppose that $i$ confirmed block ${\sf B}$ at time $t$, and at time $\tau>t+4\Delta$ node $i$ received block ${\sf B}^{\prime}\neq{\sf B}$ that conflicts with block ${\sf B}$. Then $i$ ignores block ${\sf B}^{\prime}$ and any block that extends it.
In other words, the Stubborn Nakamoto protocol differs from the standard Nakamoto protocol by (i) an explicit condition for finality, (ii) a $2\Delta$ conflict-discovery delay before finalizing a block, and (iii) a direction to miners to ignore any block that contradicts a sufficiently old finalized block. The explicit condition for finality provides an honest node with a guarantee of the behavior of other honest nodes. The $2\Delta$ delay before a node finalizes a block serves as a conflict-discovery window: it ensures that all honest nodes finalize the same block even when the attacker controls arbitrary resources.
@@
One interpretation of the protocol is that it incorporates the community response (discussed in Section 3.5) into the protocol itself. An attacker can create a conflicting ledger because the Stubborn Nakamoto protocol is permissionless. The protocol prevents the attack described in Section 3.2 by using synchrony to identify the attack: any honest nodes can observe that attack blocks arrive abnormally late. If a node can identify received blocks as attack blocks that can be mined only by an attacker, it can simply ignore them. However, to maintain agreement, we must ensure that if node $i$ ignores a block, it can be confident that any honest node $j$ will also ignore that block. More generally, the protocol ensures that all honest nodes are in agreement even when an attacker creates arbitrary conflicting blocks. The Stubborn Nakamoto protocol achieves this through the agreed-upon finality rule and conflict-discovery window.
@@
Given the impossibility of consensus without an honest majority (Pass and Shi, 2017a, ; Lewis-Pye and Roughgarden,, 2023), a consistent protocol must lose liveness under some attacks. We discuss how the protocol can regain liveness after an attack in Section 4.3.
@@ 4.2 Consistency and liveness of Stubborn Nakamoto
Under synchrony, Stubborn Nakamoto satisfies consistency even in the presence of an adversary with unbounded computational power and any number of corrupted nodes.
@@
Suppose that honest node $i$ confirms block ${\sf B}$ at time $t$ and later finalizes this block after waiting $2\Delta$. We need to show that if any honest node $j$ finalized a block ${\sf B}^{\prime}$ of the same height as ${\sf B}$, then it must be that ${\sf B}^{\prime}={\sf B}$.
It is immediate from the definition that node $i$ never finalizes a block ${\sf B}^{\prime}\neq{\sf B}$ that conflicts with ${\sf B}$. Consider an honest node $j\neq i$ that finalized a block ${\sf B}^{\prime}$ of the same height as ${\sf B}$. Observe that if some node $j$ were to confirm some block ${\sf B}^{\prime}$ at time $\tau\leq t+3\Delta$, then $i$ receives and confirmes block ${\sf B}^{\prime}$ by time $\tau+\Delta$.
For the sake of contradiction, suppose that at time $\tau$ node $j$ confirms a block ${\sf B}^{\prime}$ that conflicts with ${\sf B}$. Since $i$ did not confirm any block that conflicts with ${\sf B}$ by time $t+2\Delta$, we must have that $\tau>t+\Delta$. Because node $i$ confirms block ${\sf B}$ at time $t$, it must be that node $j$ receives and confirms block ${\sf B}$ by time $t+\Delta<\tau$. Thus, node $j$ must have halted after confirming both ${\sf B},{\sf B}^{\prime}$ and did not finalize ${\sf B}^{\prime}$, leading to a contradiction.
Thus, if node $j$ finalized a block ${\sf B}^{\prime}$ of the same height as ${\sf B}$, it must be that ${\sf B}={\sf B}^{\prime}$. ∎
Theorem 1 shows that the Stubborn Nakamoto protocol is infinitely economically secure, as even an attacker with unlimited resources cannot cause a consistency violation.
To see how Stubborn Nakamoto also satisfies liveness, observe that Stubborn Nakamoto differs from Nakamoto only when, according to the local view of some honest node $i$, there are two conflicting blocks ${\sf B}^{\prime}\neq{\sf B}$ of the same height $h({\sf B})=h({\sf B}^{\prime})$, and the two blocks were both at least $k$-deep in the longest chain in view. In other words, nodes will act differently under the two protocols only in scenarios very close to the Nakamoto protocol’s consistency violations. Under approximately the same conditions that guarantee consistency and liveness of the Nakamoto protocol, such scenarios do not arise, and Stubborn Nakamoto satisfies consistency and liveness.
Theorem 2 states that Stubborn Nakamoto satisfies consistency and liveness under a sufficient honest majority assumption, approximately the same assumption required for Nakamoto to satisfy consistency and liveness. The exact technical statement of the network-aware honest majority assumption and the proof can be found in Appendix A.
@@
Under synchrony, Stubborn Nakamoto satisfies consistency and liveness if the network-aware honest majority assumption holds. 66todo: 6Poly-length run, assumption on $k$ and $\lambda$
Roughly speaking, the theorem states that under approximately the same honest majority assumption that guarantees consistency and liveness of Nakamoto, we also have consistency and liveness of Stubborn Nakamoto. When the sufficient honest majority does not hold, Nakamoto may suffer consistency violations. Given a scenario that results in a consistency violation under Nakamoto, Stubborn Nakamoto will either maintain consistency and liveness (e.g., by having all online nodes ignore an attack chain) or suffer a liveness violation while maintaining consistency (e.g., by having nodes that observe conflicting confirmations halt).
@@
The proof of Theorem 2 uses the intermediate stochastic bounds used to prove the consistency and liveness of Nakamoto. The scenarios under which liveness of Stubborn Nakamoto is violated are very close to those under which consistency of Nakamoto is violated. However, it is possible to have a liveness violation of Stubborn Nakamoto without a consistency violation of Nakamoto. For this reason, the proof of Theorem 2 requires subtle changes to the analysis of Nakamoto consensus of Pass et al., (2017) and Shi, (2022). We state the required lemmas and prove Theorem 2 in Appendix A.
Finally, we remark that some trade-offs in protocol design are necessary. No protocol can guarantee both consistency and liveness under corrupt majority (Pass and Shi, 2017a, ; Lewis-Pye and Roughgarden,, 2023). Roughly speaking, by detecting conflicts, Stubborn Nakamoto transforms a consistency violation of Nakamoto into a liveness violation.
77todo: 7generalize to pbft style protocols

### v2

@@ 4.1 The Stubborn Nakamoto protocol
The Stubborn Nakamoto protocol requires the same communication assumptions as Bitcoin’s Nakamoto: synchronous communication with late joiners. We make the standard “implicit echoing” assumption (which is also assumed for standard Nakamoto). [^54] That is, every node (including observer nodes that do not mine) echoes every fresh message they see. This implies that if an honest node sees a message at time $t$, then every honest node will have seen it by $t+\Delta$. It also shares the same setup assumption as Bitcoin’s Nakamoto: all nodes know the genesis block and the protocol. Stubborn Nakamoto is parameterized by a confirmation depth $k$.
The Stubborn Nakamoto protocol directs nodes to take two steps before finalizing a block and adding it to their local copy of the ledger. First, the block needs to be certified. Second, the node waits to see that no other node could have certified a conflicting block. If the node does not observe any potential conflicts, the node can finalize the block and add it to the ledger.
Definition 5 (The Stubborn Nakamoto Protocol).
As in Nakamoto, the protocol directs nodes (miners) to attempt to mine a block that extends the longest chain in view. Each node $i$ takes the following steps [^55] Using notation: Let $t_{1}$ be the earliest time such that ${\sf B}\in\mathcal{C}_{i}^{t_{1}}$. 1. If ${\sf B}\not\sim\mathcal{C}_{i}^{t_{1}}$, then node $i$ halts at time $t_{1}$. 2. Let $t_{2}=t_{1}+2\Delta$. If ${\sf B}\not\sim\mathcal{PC}_{i}^{t_{2}}(2\Delta)$, then node $i$ halts at time $t_{2}$. If ${\sf B}\sim\mathcal{PC}_{i}^{t_{2}}(2\Delta)$, then ${\sf B}\in{{\sf Log}}_{i}^{t_{2}}$. 3. Let $t_{3}=t_{1}+4\Delta$. If ${\sf B}\not\sim\mathcal{PC}_{i}^{t_{3}}(4\Delta)$, then node $i$ halts at time $t_{3}$. towards finalizing blocks:
  1. 1.
    Node $i$ certifies blocks according to definition 3. If a block ${\sf B}$ is certified but is not clear-certified (that is, in the local view of node $i$ there is a block ${\sf B}^{\prime}$ that is certified and conflicts with ${\sf B}$), then node $i$ halts.
  2. 2.
    If $i$ certified block ${\sf B}$ at time $t_{1}$, then at time $t_{2}=t_{1}+2\Delta$ node $i$ checks if there is a block ${\sf B}^{\prime}$ in $i$’s local view that conflicts with ${\sf B}$ and is $2\Delta$-potentially certified. If yes, node $i$ halts. If ${\sf B}$ does not conflict with any $2\Delta$-potentially certified block in $i$’s local view, then node $i$ finalizes ${\sf B}$, written ${\sf B}\in{{\sf Log}}_{i}^{t_{2}}$.
  3. 3.
    If $i$ certified block ${\sf B}$ at time $t_{1}$, then at time $t_{3}=t_{1}+4\Delta$ node $i$ checks if there is a block ${\sf B}^{\prime}$ in $i$’s local view that conflicts with ${\sf B}$ and is $4\Delta$-potentially certified. If yes, node $i$ halts.
The definition of the Stubborn Nakamoto protocol is conceptually simple yet carefully designed. On one hand, the protocol ensures that nodes can detect and prevent conflicts with other nodes. On the other hand, nodes must not be overly cautious, as halting too easily would hinder progress and compromise the protocol’s liveness. The protocol’s design, along with the definition of $\gamma$-potentially confirmed blocks, carefully balances these competing requirements.
@@
One interpretation of the protocol is that it incorporates the community response (discussed in Section 3.5) into the protocol itself. Because the Stubborn Nakamoto protocol is permissionless, it cannot prevent an attacker with sufficient resources from creating a conflicting ledger (as described in Section 3.2). Despite this, the protocol prevents consistency violations by allowing nodes to identify attacks. If a node can identify received blocks as attack blocks that could only have been mined by an attacker, it can simply ignore them. However, to maintain agreement, we must ensure that if node $i$ ignores a block, it can be confident that any honest node $j$ will also ignore that block.
To ensure agreement among all honest nodes, the Stubborn Nakamoto protocol differs from the standard Nakamoto in two key ways. First, it specifies an explicit condition for finality, ensuring that an honest node can reliably predict the behavior of other honest nodes. Second, before finalizing a block, nodes wait for a conflict-discovery period, allowing them to learn if another node receives a conflicting ledger. While an attacker with unlimited resources may still create a conflicting ledger, honest nodes will be able to detect the existence of multiple potential ledgers, identify the risk of disagreement, and halt if necessary. More generally, the protocol guarantees that an honest node finalizes a block only if that node is assured that no other honest node could have finalized a conflicting block.
@@
The Stubborn Nakamoto protocol prioritizes consistency and instructs nodes to finalize a block only if consistency is assured. Given the known impossibility of consensus without an honest majority (Pass and Shi, 2017a, ; Lewis-Pye and Roughgarden,, 2023; Sridhar et al.,, 2024), any protocol must, under some attacks, sacrifice either consistency or liveness. The protocol may lose liveness under some attacks that cause nodes to detect a conflict and halt. We discuss how the protocol can regain liveness after an attack in Section 4.4.
Remark 1.
The Stubborn Nakamoto protocol can be made more resilient by instructing nodes to become “stubborn” on any block for which they completed step 3, meaning a node ignores any block that conflicts with a block that completed step 3 without halting. This modification allows online nodes to completely ignore the attack described in the Section 3.2, and it does not affect the protocol’s security and liveness guarantees for online nodes (which are formally stated and proved in the following section). However, such a modification will require careful treatment of late joiners, which we discuss in Section 4.3 and Section 4.4.
@@ 4.2 Consistency and liveness of Stubborn Nakamoto
We turn to prove that Stubborn Nakamoto guarantees consistency and liveness for online nodes. Treatment of late joiners is in Section 4.3. Before presenting our main theorem, we first establish formal results that characterize the knowledge an honest online node has about other honest online nodes under the Stubborn Nakamoto protocol.
Lemma 1.
If an honest node $i$ certifies block ${\sf B}$ at time $t$, then by time $t+\Delta$ block ${\sf B}$ is $2\Delta$-potentially certified by any honest node $j$.
If an honest node $i$ $\gamma$-potentially certifies block ${\sf B}$ at time $t$, then by time $t+\Delta$ block ${\sf B}$ is $(\gamma+2\Delta)$-potentially certified by any honest node $j$.
Proof.
Because $\mathcal{C}_{i}^{t}=\mathcal{PC}_{i}^{t}(0)$, it is sufficient to prove the second part of the lemma. Assume that block ${\sf B}$ is $\gamma$-potentially certified by honest node $i$ at time $t$. That is, there is a chain $\sf Chain(\bar{{\sf B}})=({\sf B}_{0},\dots,{\sf B},\dots,\bar{{\sf B}})\subseteq{\sf View}_{i}^{t}$ that satisfies the conditions of definition 4 given the local view of node $i$ at time $t$, namely: $h(\bar{{\sf B}})-h({\sf B})+1\geq k$, and $h(\bar{{\sf B}})\geq h(\tilde{{\sf B}})$ for any $\tilde{{\sf B}}\in{{\sf View}}_{i}^{t-\gamma}$
Consider arbitrary honest node $j$. Because node $i$ receives all the blocks in $\sf Chain(\bar{{\sf B}})$ by time $t$, node $j$ receives all blocks in $\sf Chain(\bar{{\sf B}})$ by time $\tau\leq t+\Delta$. We have that ${\sf B}$ is $k$-deep in $\sf Chain(\bar{{\sf B}})$, satisfying $h(\bar{{\sf B}})-h({\sf B})+1\geq k$. To see that $h(\bar{{\sf B}})\geq h(\tilde{{\sf B}})$ for any $\tilde{{\sf B}}\in{{\sf View}}_{j}^{\tau-\gamma-2\Delta}$, observe that
${{\sf View}}_{j}^{\tau-\gamma-2\Delta}\subseteq{{\sf View}}_{j}^{t+\Delta-\gamma-2\Delta}={{\sf View}}_{j}^{t-\gamma-\Delta}\subseteq{{\sf View}}_{i}^{t-\gamma}$
where the first inclusion follows because $\tau\leq t+\Delta$, and the last inclusion follows from communication between two honest nodes. Thus, we have that ${\sf B}\in\mathcal{PC}_{j}^{\tau}(2\Delta+\gamma)\subseteq\mathcal{PC}_{j}^{t+\Delta}(2\Delta+\gamma)$. ∎
We can now state our main result, showing that Stubborn Nakamoto is economically secure.
@@
Under synchrony, Stubborn Nakamoto satisfies consistency even in the presence of an adversary with unbounded computational power [^56] That is, the attacker can control arbitrarily mining resources with an unbounded hash rate. We maintain the standard assumption that the attacker cannot break cryptographic primitives. and any number of corrupted nodes.
@@
Suppose that honest node $i$ finalized block ${\sf B}$ at time $t_{2}$. Let $t_{1}=t_{2}-2\Delta$ denote the time when node $i$ certified block ${\sf B}$. We need to show that if some honest node $j$ finalized a block ${\sf B}^{\prime}$ then it must be that ${\sf B}^{\prime}\sim{\sf B}$. It is immediate from definition 3 that node $i$ never clear-certifies a block ${\sf B}^{\prime}\not\sim{\sf B}$, and therefore node $i$ cannot finalize a block ${\sf B}^{\prime}$ that conflicts with ${\sf B}$.
For the sake of contradiction, consider an honest node $j\neq i$ that finalized a block ${\sf B}^{\prime}\not\sim{\sf B}$ at time $t_{2}^{\prime}$, and let $t_{1}^{\prime}=t_{2}^{\prime}-2\Delta$ denote the time node $j$ certified block ${\sf B}^{\prime}$. By 1, node $j$ must have $2\Delta$-potentially certified block ${\sf B}$ by time $t_{1}+\Delta$. Because node $j$ finalized block ${\sf B}^{\prime}\not\sim{\sf B}$, we must have that $t_{1}^{\prime}+2\Delta=t_{2}^{\prime}<t_{1}+\Delta$. Symmetrically, by 1 we have that node $i$ must have $2\Delta$-potentially certified block ${\sf B}^{\prime}$ by time $t_{1}^{\prime}+\Delta$. Because node $i$ finalized block ${\sf B}\not\sim{\sf B}^{\prime}$, we must have that $t_{1}+2\Delta=t_{2}<t_{1}^{\prime}+\Delta$. But this implies that $t_{1}<t_{1}^{\prime}-\Delta$ and $t_{1}^{\prime}<t_{1}-\Delta$, a contradiction.
The consistency guarantee provided by Footnote 56 is notable for two key reasons. First, it demonstrates that the Stubborn Nakamoto protocol is maximally economically secure, as even an attacker with unlimited resources cannot violate consistency and deceive the protocol’s users. Second, the protocol’s security does not rely on mining rewards or other monetary incentives. [^57] Mining rewards play a role in deterring attackers from compromising the protocol’s liveness; see Section 4.4.
Of course, a protocol is only useful if it also satisfies liveness. To see why Stubborn Nakamoto also satisfies liveness, observe that Stubborn Nakamoto differs from Nakamoto only when, in the local view of some honest node $i$, there are two conflicting blocks ${\sf B}^{\prime}\not\sim{\sf B}$ that are both $\gamma$-potentially certified for $\gamma\leq 4\Delta$. In other words, nodes will act differently under the two protocols only in realizations that are close to the Nakamoto protocol’s consistency violations. Under approximately the same conditions that guarantee consistency and liveness of the Nakamoto protocol, such realizations do not arise, and Stubborn Nakamoto satisfies consistency and liveness.
Theorem 2 builds on these ideas to formally establish that Stubborn Nakamoto satisfies consistency and liveness under a sufficient honest majority assumption, the same assumption required for Nakamoto to satisfy consistency and liveness. The exact technical statement of the network-aware honest majority assumption and the proof can be found in Appendix A.
@@
Under synchrony, Stubborn Nakamoto satisfies consistency and liveness if the network-aware honest majority assumption holds.
Roughly speaking, the theorem states both Nakamoto and Stubborn Nakamoto require the same honest majority assumption to guarantee both consistency and liveness. The protocols differ when this assumption is not satisfied. [^58] Some trade-offs in protocol design are necessary, as no protocol can guarantee both consistency and liveness under corrupt majority (Pass and Shi, 2017a, ; Lewis-Pye and Roughgarden,, 2023; Sridhar et al.,, 2024). Under the Nakamoto protocol, an attack that controls sufficiently many nodes can cause consistency violations. In contrast, the Stubborn Nakamoto protocol prevents consistency violations by allowing nodes to detect when consistency might be at risk. Rather than allowing transactions to be added to the ledger only to be later reverted, Stubborn Nakamoto instructs nodes to halt and wait for a safe recovery. Effectively, the Stubborn Nakamoto protocol transforms Nakamoto’s consistency violations into liveness violations. This behavior is particularly desirable in financial systems, where transaction integrity is paramount.
Note that the version of Stubborn Nakamoto outlined in Remark 1 can completely avoid some of Nakamoto’s consistency violations without suffering a liveness violation. In particular, this version of the protocol maintains both consistency and liveness under the attack outlined in Section 3.2.
@@
The proof of Theorem 2 uses the intermediate stochastic bounds used to prove the consistency and liveness of Nakamoto. While the scenarios under which liveness of Stubborn Nakamoto is violated are very close to those under which consistency of Nakamoto is violated, it is possible to have a liveness violation of Stubborn Nakamoto without a corresponding consistency violation of Nakamoto. For this reason, the proof of Theorem 2 requires subtle changes to the analysis of Nakamoto consensus of Pass et al., (2017) and Shi, (2022). We state the technical details, required lemmas, and the proof of Theorem 2 in Appendix A.
4.3 Late joiners
This section establishes guarantees for late joiners under Stubborn Nakamoto. A late joiner is a node that joins the system at some time $t$, after the protocol has already been running for some time and some blocks may already have been finalized. We adopt the standard communication assumption that by time $t+\Delta$ the late joiner receives all messages sent by honest nodes by time $t$ or earlier (Pass and Shi, 2017b, ).595959 The same communication assumption is also used by the standard Nakamoto protocol, which requires late-joining nodes to receive and verify the entire chain of previous blocks, linking back to the genesis block. The delivery of past messages is essential in any protocol to ensure that newly joining nodes can correctly establish and continue the ledger. However, a late joiner node cannot distinguish the time at which earlier messages were sent. [^60] That is, if $\tau_{1},\tau_{2}<t$, a late joiner cannot tell whether a message was sent at time $\tau_{1}$ or at time $\tau_{2}$. This is in contrast to online nodes, which are able to observe when any message is sent by an honest node up to a $\Delta$ time lag. In particular, late joiners cannot recreate the finalization steps taken by online nodes under Stubborn Nakamoto, as they cannot determine which blocks were received within $4\Delta$ of a block’s certification. Additionally, a late joiner may also receive arbitrarily many messages from adversarial nodes.
Allowing late joiners is crucial for the free entry of nodes. Without this capability, all nodes would be required to join at the protocol’s inception and remain continuously online. However, ensuring consistency for late joiners presents significant challenges. For example, the well-known Dolev-Strong protocol guarantees consistency and liveness for online nodes but fails to provide these guarantees for late joiners (see, e.g., Pass and Shi, 2017a ).
The following definition formalizes the process by which a late joiner node becomes an online node. We refer to these steps as the node’s wake-up procedure.
Definition 6 (Node Awakening).
Let honest node $i$ be a late joiner that awakens at time $t_{w}$. Node $i$ waits until time $t_{on}>t_{w}$ such that node $i$ does not receive any new blocks between $t_{on}-6\Delta$ and $t_{on}$. If there are two conflicting blocks ${\sf B},{\sf B}^{\prime}\in{\sf View}_{i}^{t_{on}}$ that are both $k$-deep in some longest chain in ${\sf View}_{i}^{t_{on}}$, then node $i$ halts. Otherwise, node $i$ certifies and finalizes all blocks that are $k$-deep in a longest chain in view, and becomes an online honest node.
The following result establishes the protocol’s consistency guarantee for late joiners. The Stubborn Nakamoto protocol requires an additional assumption to ensure consistency for late joiners: namely, a late joiner must rely on at least one online honest node to maintain consistency with blocks that were finalized before the late joiner woke up.
Theorem 3.
Suppose that node $i$ is a late joiner that wakes up according to Definition 6, and there is some honest node $j$ that is online throughout and did not halt by time $t_{on}$. Then consistency holds for node $i$ even in the presence of an adversary with unbounded computational power and any number of corrupted nodes.
Remark 2.
Theorem 3 does not hold under the protocol modification outlined in Remark 1. The reason is that while online nodes can detect and ignore the attack chain described in Section 3.2, late joiners lack the ability to distinguish between the attack chain and the honest chain. In other words, late joiners do not have sufficient information to select the correct chain after online nodes ignore an attack chain. These challenges are addressed by the recovery oracle, introduced in Section 4.4.
Proof.
Suppose that $i$ is a late joiner, and became online at time $t_{on}$. Let $j$ be an honest node that is online and did not halt by time $t_{on}$.
First, let us show that ${\sf Log}^{t_{on}}_{i}={\sf Log}^{t_{on}}_{j}$. Let ${\sf B}_{j}\in{\sf Log}^{t_{on}}_{j}$ be a block finalized by $j$. Let $t_{j}$ denote the time at which block ${\sf B}_{j}$ was certified by $j$ and let $\sf Chain(\bar{{\sf B}}_{j})$ denote the corresponding chain that certified ${\sf B}_{j}$. Because ${\sf B}_{j}$ is finalized by $j$ by time ${t_{on}}$, we have that $t_{j}\leq{t_{on}}-2\Delta$. Therefore, $\sf Chain(\bar{{\sf B}}_{j})\subseteq{\sf View}_{i}^{t_{on}}$. If a longest chain in ${\sf View}_{i}^{t_{on}}$ contains ${\sf B}_{j}$, it must be as at least as long as $\sf Chain(\bar{{\sf B}}_{j})\subseteq{\sf View}_{i}^{t_{on}}$, and therefore include ${\sf B}_{j}$ at least $k$-deep. This implies that ${\sf B}_{j}\in{\sf Log}^{t_{on}}_{i}$, as required.
Otherwise, assume for the sake of contradiction that $\sf Chain(\bar{{\sf B}}^{\prime})$ is a longest chain in ${\sf View}_{i}^{t_{on}}$ that does not include ${\sf B}_{j}$. Since $\sf Chain(\bar{{\sf B}}^{\prime})$ is at least as long as $\sf Chain(\bar{{\sf B}}_{j})$, it must include some block ${\sf B}^{\prime}\not\sim{\sf B}$ that is at least $k$-deep in $\sf Chain(\bar{{\sf B}}^{\prime})$. We have that
${\sf View}_{i}^{{t_{on}}-6\Delta}\subseteq{\sf View}_{j}^{{t_{on}}-5\Delta}\subseteq{\sf View}_{j}^{{t_{on}}-\Delta}\subseteq{\sf View}_{i}^{{t_{on}}}\;,$
and therefore $\sf Chain(\bar{{\sf B}}^{\prime})$ is a longest chain in $j$’s view at time ${t_{on}}-5\Delta$. But this implies that node $j$ must halt before time ${t_{on}}$, having seen a conflicting certified block.
To show ${\sf Log}^{t_{on}}_{i}\subseteq{\sf Log}^{t_{on}}_{j}$, let ${\sf Log}^{t_{on}}_{i}=\sf Chain({\sf B}_{i})$ and let $\sf Chain(\bar{{\sf B}}_{i})$ denote a chain that certified ${\sf B}_{i}$. We have that $h(\bar{{\sf B}}_{i})\geq h(\tilde{B})$ for all $\tilde{{\sf B}}\in{\sf View}_{i}^{{t_{on}}}$. Moreover, because $i$ does not halt by time ${t_{on}}$, ${\sf B}_{i}$ is $k$ deep in any longest chain in ${\sf View}_{i}^{{t_{on}}}$. We have that
$\sf Chain(\bar{{\sf B}}_{i})\subseteq{\sf View}_{i}^{{t_{on}}}={\sf View}_{i}^{{t_{on}}-6\Delta}\subseteq{\sf View}_{j}^{{t_{on}}-5\Delta}\;.$
Therefore, node $j$ certified ${\sf B}_{i}$ by time ${t_{on}}-5\Delta$. Because $j$ did not halt by time ${t_{on}}$, we have that $j$ finalized ${\sf B}_{i}$ and every block in $\sf Chain({\sf B}_{i})$ by time ${t_{on}}$. Thus, we have that ${\sf Log}^{t_{on}}_{i}={\sf Log}^{t_{on}}_{j}$.
Now, suppose that node $i$ certifies block ${\sf B}$ at time $t_{1}>{t_{on}}$ and finalizes it at time $t_{2}=t_{1}+2\Delta$. It is immediate that ${\sf B}$ does not conflict with any block certified or finalized by $i$. Suppose, for the sake of contradiction, that honest node $j$ finalizes a conflicting block ${\sf B}^{\prime}$ at time $t_{2}^{\prime}$. If $t_{2}^{\prime}\leq{t_{on}}$, we have ${\sf B}^{\prime}\in{\sf Log}^{t_{on}}_{j}={\sf Log}^{t_{on}}_{i}$ by the first part of the proof, and we have a contradiction.
If $t_{2}^{\prime}>{t_{on}}$, it must be that $j$ certifies the block ${\sf B}^{\prime}$ at time $t_{1}^{\prime}>{t_{on}}-\Delta$. Let $\sf Chain(\bar{{\sf B}}^{\prime})$ be the corresponding chain in $j$’s view at time $t_{1}^{\prime}$. We have that $\sf Chain(\bar{{\sf B}}^{\prime})$ must be longer than any chain in $j$’s view at time $t_{1}^{\prime}$, and therefore it is longer than any chain in ${\sf View}_{i}^{{t_{on}}}$. Node $i$ receives $\sf Chain(\bar{{\sf B}}^{\prime})$ by time $\tau\leq t_{1}^{\prime}+\Delta$. Because ${\sf View}_{i}^{\tau-2\Delta}\subseteq{\sf View}_{j}^{\tau-\Delta}\subseteq{\sf View}_{j}^{t_{1}^{\prime}}$, we have that ${\sf B}^{\prime}\in\mathcal{PC}_{i}^{\tau}(2\Delta)$. Since node $i$ finalized block ${\sf B}$, we must have $\tau>t_{2}$.
Let $\sf Chain(\bar{{\sf B}})$ be the chain in $i$ view that certified ${\sf B}$ at time $t_{1}$. Again, we have that $\sf Chain(\bar{{\sf B}})$ is longer than any chain in ${\sf View}_{i}^{{t_{on}}}$. Node $j$ receives $\sf Chain(\bar{{\sf B}})$ by time $\nu\leq t_{1}+\Delta$. Because ${\sf View}_{j}^{\nu-2\Delta}\subseteq{\sf View}_{i}^{\nu-\Delta}\subseteq{\sf View}_{i}^{t_{1}}$, we have that ${\sf B}\in\mathcal{PC}_{j}^{\nu}(2\Delta)$. But this implies that $t_{1}+\Delta\geq\nu>t_{1}^{\prime}+2\Delta$ and $t_{1}^{\prime}+\Delta\geq\tau>t_{1}+2\Delta$, and we have reached a contradiction.
When the honest majority assumption holds, late joiners can reliably identify previously finalized blocks that must be in the longest chain in their view. The following theorem provides the protocol’s guarantees for late joiners in the absence of an attack.
Theorem 4.
If the honest majority assumption holds, Stubborn Nakamoto satisfies consistency and liveness for online nodes and late joiners.
Proof.
The proof follows from Theorem 2 together with the observation that if consistency and liveness hold for online nodes, then Stubborn Nakamoto and the standard Nakamoto protocol generate identical executions for late joiners (in the sense that any honest node maintains the same finalized ledger). ∎